# GradMix: Gradient-based Selective Mixup for Robust Data Augmentation in Class-Incremental Learning

### v1

@@ Appendix B Appendix – Algorithm
Appendix C Appendix – Experiments
C.1 More Details on Experimental Settings
@@
C.2 More Details on Catastrophic Forgetting in Mixup
@@
Figure 4: Comparison between the performance of ER, Mixup, and GradMix on the MNIST dataset.
@@
Figure 5: Comparison between the performance of ER, Mixup, and GradMix on the Fashion-MNIST dataset.
@@
Figure 6: Comparison between the performance of ER, Mixup, and GradMix on the CIFAR-10 dataset.
@@
Figure 7: Comparison between the performance of ER, Mixup, and GradMix on the CIFAR-100 dataset.
@@
Figure 8: Comparison between the performance of ER, Mixup, and GradMix on the Tiny ImageNet dataset.
@@
Table 6: Average rate of detrimental mixed samples on catastrophic forgetting when applying Mixup to the experience replay method (ER) on the MNIST, Fashion-MNIST (FMNIST), CIFAR-10, CIFAR-100, and Tiny ImageNet datasets.
@@
C.3 More Details on the Error when using Approximated Gradients of Mixed Samples
Continuing from Sec. 5.2, we present approximation errors in inner products of the gradients of buffer data and mixed training data as shown in Fig. 9. We compute both the approximated and true inner products when using the corresponding approximated and true gradients for the mixed samples and compare them using Root Mean Squared Error (RMSE) as a metric. A lower RMSE value means better approximation. If $\lambda$ is 0 or 1, the approximation error is 0 since the approximated gradient becomes the same as the gradient of the original sample. As the $\lambda$ value gets closer to 0.5, mixed samples are generated in the middle of original samples and the resulting approximation error increases. If $\lambda$ follows a uniform distribution between 0 and 1, the average approximation errors are 0.136, 0.110, 0.074, 0.035, and 0.037 for the MNIST, FMNIST, CIFAR-10, CIFAR-100, and Tiny ImageNet datasets, respectively.
@@ C.3 More Details on the Error when using Approximated Gradients of Mixed Samples
Figure 9: Root Mean Squared Error (RMSE) between approximated and true inner products with respect to mixing ratio $\lambda$ on the MNIST, FMNIST, CIFAR-10, CIFAR-100, and Tiny ImageNet datasets.
@@
Table 7: Average precision and recall results for determining the optimal sample pairs using approximated gradients of mixed samples on the MNIST, FMNIST, CIFAR-10, CIFAR-100, and Tiny ImageNet datasets.
@@
C.4 More Details on Class-based Selective Mixup
Continuing from Sec. 5.3, we provide empirical results to show that training samples within the same class share similar gradients as shown in Fig. 10. We use the cosine distance as a metric and define intra-class and inter-class distances as the average distances between gradients of samples within the same class and from different classes, respectively. As a result, intra-class distances are smaller than inter-class distances across all datasets, indicating that training samples within the same class share similar gradients.
Figure 10: Average cosine distances between gradients of data within the same class (intra-class) and from different classes (inter-class) on the MNIST, FMNIST, CIFAR-10, CIFAR-100, and Tiny ImageNet datasets.
C.5 More Details on Computational Complexity and Runtime Results of GradMix
Continuing from Sec. 5.3, we provide more details on an analysis of the computational complexity and runtime of the GradMix algorithm. The computational complexity of GradMix is quadratic to the number of classes when solving the optimization problem to obtain the optimal class pairs for each class. However, since we transform the data selection optimization problem into a class selection problem, the overhead is significantly reduced as the number of classes is much smaller than the number of samples. If the number of classes keeps increasing with consecutive tasks, the overhead of GradMix may also increase correspondingly. If there are too many classes, we can transform the repeated gradient inner product process into a nearest neighbor search problem and use locality-sensitive hashing (LSH) [28, 23] to efficiently find an approximate solution with a time complexity of $\mathcal{O}(C\log C)$, where $C$ is the number of classes.
We also provide runtime results for solving the optimization problem in GradMix per epoch with respect to the increasing number of tasks as shown in Fig. 11. Since GradMix is not applied to the first task, we present the results starting from the second task. As the number of tasks increases, the number of classes also increases, which leads to a gradual increase in the runtime. In practice, solving the optimization problem per epoch takes a few seconds, resulting in a total overhead of a few minutes during training on the MNIST, FMNIST, CIFAR-10, CIFAR-100, and Tiny ImageNet datasets.

### v2

@@ 5.2. Gradients of Mixed Samples
5.3. Algorithm
We now design the GradMix algorithm to solve the data selection optimization problem. By substituting Eq. 6 and Eq. 7 into Eq. 5 and using the last layer approximation, we can efficiently compute the inner products of the average gradient of buffer data and the gradients of all mixed samples to find the optimal mixing sample. However, formulating a sample-based criterion remains practically infeasible, as it would require explicit per-sample gradient evaluations over the entire set of mixed samples, resulting in excessive computational and memory overhead. For better efficiency, we group samples that have similar gradients and reformulate the optimization problem from data selection to group selection by replacing the gradient of a mixed sample in Eq. 5 with the average gradient of mixed samples from two arbitrary groups. By leveraging the observation that training samples within the same class share similar gradients as described in Sec. B.4, we group samples by class and propose a class-based selective mixup criterion. If the inner product of the average gradients of buffer data and mixed samples from two classes is negative, it implies that training with the mixed samples increases the average loss of previous tasks and worsens catastrophic forgetting. In this case, we define the class pair as a negative class pair and replace the mixed samples with those generated from the optimal class pair that maximizes the inner product. Conversely, if the inner product is positive, indicating that the average loss of previous tasks does not increase, and catastrophic forgetting does not occur, we retain the mixed samples as part of the training data. Since generating mixed samples solely from the optimal class pair may reduce sample diversity from a data augmentation perspective, our design choice balances the trade-off between mitigating forgetting and enhancing diversity.
We describe the overall process of GradMix in Algorithm 1. At each epoch during training, we compute the average gradient of buffer data (denoted as $\overline{g}_{buf}$) using Eq. 6 (Step 3). Next, we set a mixing ratio $\lambda$ and compute the average gradients of mixed samples from all class pairs (denoted as $\tilde{g}_{y_{i}y_{j}}$) using Eq. 7 (Steps 4–7). By computing the inner products of $\overline{g}_{buf}$ and $\tilde{g}_{y_{i}y_{j}}$ for all class pairs, we get the set of negative class pairs (Step 8) and the set of optimal class pairs for each class (Step 9). The selective mixup criterion is determined using the two sets once per epoch and applied to all batches to ensure computational efficiency. For each batch consisting of both current task data and buffer data (Steps 11–13), we first randomly shuffle the batch to generate sample pairs to be mixed and set a mixing ratio $\lambda$ following the configuration in Mixup (Zhang et al., 2018) (Steps 14–15). However, if the class pairs of the sample pairs belong to the negative set, we replace them with sample pairs from the optimal class pairs (Steps 16–19). After this replacement process, we form a new batch consisting of mixed samples that satisfy our selective mixup criterion (Step 22) and update the model (Step 23). Lastly, we store a random subset of the current task data to the buffer before moving onto the next task (Steps 26–27). We provide an analysis of the computational complexity and runtime of GradMix algorithm in Sec. B.5, which shows that GradMix is scalable by only computing the last layer gradients of a small exemplar set for each class.
6. Experiments
Metrics
Datasets
We use a total of five datasets for the experiments as shown in Table 1: MNIST (LeCun et al., 1998), Fashion-MNIST (FMNIST) (Xiao et al., 2017), CIFAR-10 (Krizhevsky, 2009), CIFAR-100 (Krizhevsky, 2009), and Tiny ImageNet (Russakovsky et al., 2015). To setup experiments for class-incremental learning, we divide the datasets into sequences of tasks consisting of equal numbers of disjoint classes and assume that the task boundaries are available (van de Ven and Tolias, 2019; Mirzadeh et al., 2020).
Models
For MNIST and FMNIST, we train a two-layer MLP each with 256 neurons using the SGD optimizer with an initial learning rate of 0.01 for 20 epochs. For CIFAR-10, CIFAR-100, and Tiny ImageNet, we train a ResNet-18 (He et al., 2016a, b) model using the SGD optimizer with an initial learning rate of 0.1 for 50, 250, and 100 epochs, respectively. During training, we reduce the learning rate by one-tenth at 100, 150, and 200 epochs for CIFAR-100, and at 30, 60, and 90 epochs for Tiny ImageNet. The last layer of the model is shared for all the tasks for a single-head evaluation (Farquhar and Gal, 2018; Chaudhry et al., 2018). We train all models using cross-entropy loss.
Baselines
Following the experimental setups of (Kumari et al., 2022; Zhang et al., 2022b), we use ER (Chaudhry et al., 2019b), MIR (Aljundi et al., 2019a), GSS (Aljundi et al., 2019b), DER (Buzzega et al., 2020), FOSTER (Wang et al., 2022), and MEMO (Zhou et al., 2023) as basis methods for class-incremental learning and apply data augmentation techniques to them. The experiments also can be extended to include recent class-incremental learning methods (Yan et al., 2021; Chen and Chang, 2023; Zheng et al., 2025). We compare GradMix with three types of data augmentation baselines:
  • Mixup-based methods: Mixup(Zhang et al., 2018) generates synthetic training data through linear interpolations of pairs of input features and their corresponding labels; Manifold Mixup(Verma et al., 2019) performs linear interpolations of hidden layer representations and corresponding labels of training data; CutMix(Yun et al., 2019) generates new training data by randomly cutting and pasting patches between input images, while adjusting the corresponding labels based on the proportion of each patch.
  • Imbalance-aware mixup methods: Remix(Chou et al., 2020) mixes features like Mixup, but assigns labels by providing higher weights to the minority class; Balanced-MixUp(Galdran et al., 2021) performs both regular and balanced sampling when selecting pairs of training data for Mixup.
  • Policy-based method: RandAugment(Cubuk et al., 2020) randomly applies different image transformations to training data with adjustable magnitude.
Parameters
We store 32 samples per class for buffering and set the batch size to 64 for all experiments. For the hyperparameter $\alpha$ used in the beta distribution, we follow the setting of Mixup (Zhang et al., 2018). Since the Mixup paper does not specify $\alpha$ values for the MNIST and FMNIST datasets, we set $\alpha$ to 1.0, which corresponds to a uniform distribution. In practice, the performance of GradMix is not sensitive to the value of $\alpha$ as shown in Sec. B.6. See Sec. B.1 for more experimental details.
Table 1. For each of the five datasets, we show the total training set size, the number of features, the number of classes, and the number of tasks.
Dataset Size #Features #Classes #Tasks
MNIST 50K $\times28$ 10 5
FMNIST 60K $\times28$ 10 5
[…]
Table 2. Average accuracy results on the MNIST, FMNIST, CIFAR-10, CIFAR-100, and Tiny ImageNet datasets. We use ER as a basis method of class-incremental learning and compare GradMix with three types of data augmentation baselines: (1) mixup-based methods: Mixup, Manifold Mixup, and CutMix; (2) imbalance-aware mixup methods: Remix and Balanced-MixUp; and (3) a policy-based method: RandAugment.
Type Method MNIST FMNIST CIFAR-10 CIFAR-100 Tiny ImageNet
[…]
6.1. Model Performance Results
We compare GradMix against other data augmentation baselines on the five datasets with respect to average accuracy as shown in Table 2 and Sec. B.7. Overall, GradMix consistently achieves the best accuracy results compared to the baselines for all the datasets by considering catastrophic forgetting while augmenting data. In contrast, since existing data augmentation techniques are designed for static and balanced data, they have limitations in class-incremental learning setups with dynamic and imbalanced data. We also present sequential accuracy results for each task as shown in Fig. 3 and Sec. B.8 where GradMix shows better accuracy performance than other baselines at each task. In addition, GradMix helps further improvement of advanced class-incremental learning methods, including MIR, GSS, DER, FOSTER, and MEMO as shown in Table 6.1.
(a) MNIST.
(b) FMNIST.
(c) CIFAR-10.
(d) CIFAR-100.
Figure 3. Sequential accuracy results of GradMix and baselines on the MNIST, FMNIST, CIFAR-10, and CIFAR-100 datasets as new tasks are learned.
Table 3. Average accuracy results on the CIFAR-10, CIFAR-100, and Tiny ImageNet datasets. We use MIR, GSS, DER, FOSTER, and MEMO as basis methods of class-incremental learning and integrate GradMix.
Method CIFAR-10 CIFAR-100 Tiny ImageNet
[…]
Among the mixup-based methods, Manifold Mixup consistently shows robust results. However, another mixup-based method, CutMix, shows the worst performance on CIFAR-10, CIFAR-100, and Tiny ImageNet as it significantly reduces accuracy on previous tasks. We believe that cutting images into patches in the limited previous task data results in significant information loss. The two imbalance-aware mixup methods, Remix and Balanced-MixUp, achieve better results than Mixup as they address the data imbalance between previous and current tasks using disentangled mixing and balanced sampling techniques, respectively. The policy-based method, RandAugment, also shows underperforming results. For MNIST and FMNIST, elastic distortions across scale, position, and orientation are known to be effective, but other transformations in RandAugment like flipping and random cropping may have negative effects. For CIFAR-10, CIFAR-100, and Tiny ImageNet, the transformation technique in RandAugment that removes a region of an image is harmful to retaining the knowledge of the limited buffer data.
Considering the three types of data augmentation baselines, the two imbalance-aware mixup methods, Remix and Balanced-MixUp, show consistently solid performance against other baselines by more effectively exploiting the limited buffer data for previous tasks. However, the random selection of sample pairs for mixup leads to greater catastrophic forgetting of previous tasks and limits the improvement in average accuracy. In comparison, GradMix performs gradient-based selective mixup to minimize catastrophic forgetting of previous tasks, which leads to significantly better model accuracy results for all datasets. We also show that GradMix is effective on large-scale datasets (e.g., ImageNet-1K (Deng et al., 2009)) and variations in backbone models (e.g., ViT (Dosovitskiy et al., 2021)) in Sec. B.9 and Sec. B.10, respectively.
(a) MNIST.
(b) FMNIST.
(c) CIFAR-10.
(d) CIFAR-100.
Figure 4. Selective mixup results of GradMix on the last task for the MNIST, FMNIST, CIFAR-10, and CIFAR-100 datasets.
Table 4. Ablation study on the MNIST, FMNIST, CIFAR-10, CIFAR-100, and Tiny ImageNet datasets with two ablation scenarios when randomly selected sample pairs worsen catastrophic forgetting in Algorithm 1 (Steps 17–19): (1) using original samples without mixup; and (2) retaining random mixup. We use ER as a basis method.
  Method   MNIST   FMNIST   CIFAR-10   CIFAR-100   Tiny ImageNet
[…]
Table 5. Average accuracy results on the MNIST, FMNIST, CIFAR-10, CIFAR-100, and Tiny ImageNet datasets with respect to the buffer size per class. We use ER as a basis method of class-incremental learning and compare GradMix with various data augmentation baselines.
Method MNIST FMNIST CIFAR-10 CIFAR-100 Tiny ImageNet
NN=16 NN=64 NN=16 NN=64 NN=16 NN=64 NN=16 NN=64 NN=16 NN=64
[…]
6.2. Selective Mixup Analysis
We next analyze how GradMix selectively mixes different pairs of classes. We provide selective mixup results on the last task of the MNIST, FMNIST, CIFAR-10, and CIFAR-100 datasets as shown in Fig. 4, and the results for the Tiny ImageNet dataset are shown in Sec. B.11. Since the selective mixup results for other tasks are similar, we show the results for the last task without loss of generality. As the selective mixup criterion changes each epoch based on gradients, we sum the total number of mixup operations for each class pair combination, denoted as $(C_{i},C_{j})$, during training and present the results as a heatmap.
6.3. Ablation Study
6.4. Varying the Buffer Size
We compare the performance of GradMix with the baselines when varying the number of buffer samples per class, denoted as $N$, from 32 to 16 and 64 as shown in Table 5. We observe that GradMix shows the best performance for all datasets regardless of the buffer size. In addition, as we store more buffer samples and use them for training, the average accuracy of all the baselines and GradMix increases. For the results on the Tiny ImageNet dataset with a buffer size of 16 samples per class, we believe this buffer size is too small to train the model sufficiently, as none of the data augmentation methods improve accuracy compared to ER.
7. Conclusion
Acknowledgements.
This work was supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No. RS-2022-II220157, Robust, Fair, Extensible Data-Centric Continual Learning). This work was supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No. RS-2024-00444862, Non-invasive near-infrared based AI technology for the diagnosis and treatment of brain diseases).
Appendix A Appendix – Theory
A.1. Theoretical Analysis of Applying Mixup in Class-Incremental Learning
Theorem.
Proof.
(8) $L_{\text{ER}}$ $=\frac{1}{|\mathbb{Y}_{p}|}\sum_{y\in\mathbb{Y}_{p}}\tilde{\ell}(f_{\theta},G_{y})$
(9) $L_{\text{Mixup}}=\frac{1}{|\mathbb{Y}_{p}|}\sum_{y\in\mathbb{Y}_{p}}\ell(f_{\theta}^{l-1},G_{y})-\eta\frac{1}{|\mathbb{Y}_{p}|}\sum_{y\in\mathbb{Y}_{p}}\nabla_{\theta}\ell(f_{\theta}^{l-1},G_{y})^{\top}\nabla_{\theta}\ell(f_{\theta}^{l-1},\tilde{d}_{ij}).$
By subtracting Eq. 9 from Eq. 8,
(10) $L_{\text{ER}}-L_{\text{Mixup}}$ $=-\eta\frac{1}{|\mathbb{Y}_{p}|}\sum_{y\in\mathbb{Y}_{p}}\Bigl(\nabla_{\theta}\ell(f_{\theta}^{l-1},G_{y})^{\top}\nabla_{\theta}\ell(f_{\theta}^{l-1},d_{i})-\nabla_{\theta}\ell(f_{\theta}^{l-1},G_{y})^{\top}\nabla_{\theta}\ell(f_{\theta}^{l-1},\tilde{d}_{ij})\Bigr)$
$L_{\text{ER}}-L_{\text{Mixup}}<0.$
A.2. Theoretical Analysis of the Gradients of Mixed Samples
Lemma.
Proof.
$\hat{y}=\sigma(z)=\sigma(w\cdot X+b).$
(11) $\frac{\partial\ell}{\partial z_{j}}=-\frac{\partial}{\partial z_{j}}\sum_{i=1}^{c}y_{i}\cdot\log(\hat{y}_{i})=-\sum_{i=1}^{c}y_{i}\cdot\frac{\partial}{\partial z_{j}}\log(\hat{y}_{i}).$
(12) $\frac{\partial}{\partial z_{j}}\log(\hat{y}_{i})$ $=\frac{\partial}{\partial z_{j}}\log\biggl(\frac{e^{z_{i}}}{\sum_{n=1}^{c}e^{z_{n}}}\biggr)$
$=\mathbbm{1}\{i=j\}-\frac{e^{z_{j}}}{\sum_{n=1}^{c}e^{z_{n}}}$
$=\mathbbm{1}\{i=j\}-\hat{y}_{j},$
(13) $\frac{\partial\ell}{\partial z_{j}}$ $=-\sum_{i=1}^{c}y_{i}\cdot(\mathbbm{1}\{i=j\}-\hat{y}_{j})$
$=\sum_{i=1}^{c}y_{i}\cdot\hat{y}_{j}-\sum_{i=1}^{c}y_{i}\cdot\mathbbm{1}\{i=j\}$
$\frac{\partial\ell}{\partial z_{j}}=\hat{y}_{j}-y_{j}.$
In the vector notation, we get:
$\frac{\partial\ell}{\partial z}=\hat{y}-y.$
(14) $g=(\nabla_{b}\ell(f_{\theta}^{l-1},d),\nabla_{w}\ell(f_{\theta}^{l-1},d))^{\top}=(\hat{y}-y,(\hat{y}-y)^{\top}X)^{\top}.$
Theorem.
Proof.
(15) $\tilde{g}_{ij}$ $=(\hat{y}_{ij}-\tilde{y}_{ij},(\hat{y}_{ij}-\tilde{y}_{ij})^{\top}X_{ij})^{\top}$
$\hat{y}=\sigma(f_{\theta}^{l-1}(x))=\sigma(\theta^{\top}x).$
Then, the logits for $x_{i}$ and $x_{j}$ become:
$z_{ij}$ $=\theta^{\top}\tilde{x}_{ij}$
$=\theta^{\top}(\lambda x_{i}+(1-\lambda)x_{j})$
$=\lambda\theta^{\top}x_{i}+(1-\lambda)\theta^{\top}x_{j}$
(16) $\hat{y}_{ij}$ $=\sigma(z_{ij})$
(17) $X_{ij}=\lambda X_{i}+(1-\lambda)X_{j}.$
Appendix B Appendix – Experiments
B.1. More Details on Experimental Settings
@@
B.2. More Details on Catastrophic Forgetting in Mixup
@@
Figure 5. Comparison between the performance of ER, Mixup, and GradMix on the MNIST dataset.
@@
Figure 6. Comparison between the performance of ER, Mixup, and GradMix on the Fashion-MNIST dataset.
@@
Figure 7. Comparison between the performance of ER, Mixup, and GradMix on the CIFAR-10 dataset.
@@
Figure 8. Comparison between the performance of ER, Mixup, and GradMix on the CIFAR-100 dataset.
@@
Figure 9. Comparison between the performance of ER, Mixup, and GradMix on the Tiny ImageNet dataset.
@@
Table 6. Average rate of detrimental mixed samples on catastrophic forgetting when applying Mixup to the experience replay method (ER) on the MNIST, Fashion-MNIST (FMNIST), CIFAR-10, CIFAR-100, and Tiny ImageNet datasets.
@@
B.3. More Details on the Error when using Approximated Gradients of Mixed Samples
Continuing from Sec. 5.2, we present approximation errors in inner products of the gradients of buffer data and mixed training data as shown in Fig. 10. We compute both the approximated and true inner products when using the corresponding approximated and true gradients for the mixed samples and compare them using Root Mean Squared Error (RMSE) as a metric. A lower RMSE value means better approximation. If $\lambda$ is 0 or 1, the approximation error is 0 since the approximated gradient becomes the same as the gradient of the original sample. As the $\lambda$ value gets closer to 0.5, mixed samples are generated in the middle of original samples and the resulting approximation error increases. If $\lambda$ follows a uniform distribution between 0 and 1, the average approximation errors are 0.136, 0.110, 0.074, 0.035, and 0.037 for the MNIST, FMNIST, CIFAR-10, CIFAR-100, and Tiny ImageNet datasets, respectively. We note that the actual effect of the approximation errors on the selective mixup criterion is negligible because the sign of inner products is a key measure in GradMix rather than their exact values.
@@ B.3. More Details on the Error when using Approximated Gradients of Mixed Samples
Figure 10. Root Mean Squared Error (RMSE) between approximated and true inner products with respect to mixing ratio $\lambda$ on the MNIST, FMNIST, CIFAR-10, CIFAR-100, and Tiny ImageNet datasets.
@@
Table 7. Average precision and recall results for determining the optimal sample pairs using approximated gradients of mixed samples on the MNIST, FMNIST, CIFAR-10, CIFAR-100, and Tiny ImageNet datasets.
@@
B.4. More Details on Class-based Selective Mixup
Continuing from Sec. 5.3, we provide empirical results to show that training samples within the same class share similar gradients as shown in Fig. 11. We use the cosine distance as a metric and define intra-class and inter-class distances as the average distances between gradients of samples within the same class and from different classes, respectively. As a result, intra-class distances are smaller than inter-class distances across all datasets, indicating that training samples within the same class share similar gradients. We believe that a difference between intra-class and inter-class distances still exists on large datasets because intra-class gradients are derived from similar embeddings within the same class, unlike the inter-class case.
Figure 11. Average cosine distances between gradients of data within the same class (intra-class) and from different classes (inter-class) on the MNIST, FMNIST, CIFAR-10, CIFAR-100, and Tiny ImageNet datasets.
B.5. More Details on Computational Complexity and Runtime Results of GradMix
Continuing from Sec. 5.3, we provide more details on an analysis of the computational complexity and runtime of the GradMix algorithm. The computational complexity of GradMix is quadratic to the number of classes when solving the optimization problem to obtain the optimal class pairs for each class. However, since we transform the data selection optimization problem into a class selection problem, the overhead is significantly reduced as the number of classes is much smaller than the number of samples. If the number of classes keeps increasing with consecutive tasks, the overhead of GradMix may also increase correspondingly. If there are too many classes, we can transform the repeated gradient inner product process into a nearest neighbor search problem and use locality-sensitive hashing (LSH) (Indyk and Motwani, 1998; Gionis et al., 1999) to efficiently find an approximate solution with a time complexity of $\mathcal{O}(C\log C)$, where $C$ is the number of classes.
We also provide runtime results for solving the optimization problem in GradMix per epoch with respect to the increasing number of tasks as shown in Fig. 12. Since GradMix is not applied to the first task, we present the results starting from the second task. As the number of tasks increases, the number of classes also increases, which leads to a gradual increase in the runtime. In practice, solving the optimization problem per epoch takes a few seconds, resulting in a total overhead of a few minutes during training on the MNIST, FMNIST, CIFAR-10, CIFAR-100, and Tiny ImageNet datasets. If the number of classes becomes too large, we suggest clustering similar classes and assigning the selective mixup criterion based on clusters rather than classes as a potential solution to reduce the computational overhead.